Solve for x.
P(x-1)+7= r(x-4)

Answer by nerdybill(7384)

P(x-1)+7= r(x-4)
Px-P+7= xr-4r
Px-xr-P+7= -4r
Px-xr+7= P-4r
Px-xr= P-4r-7
x(P-r) = (P-4r-7)
x = (P-4r-7)/(P-r)